# Privacy risk in GeoData: A survey

### v1

@@ 3.1.4.2 Spatial tessellation-based techniques
    Voronoi masking: Voronoi polygons define areas where the boundaries are equidistant between the surrounding points, or where inside the polygons is closer to the corresponding point than to any other point (Seidl et al., 2015). In voronoi masking, each point is snapped to the closest part along the edges of its corresponding Voronoi polygon. An advantage of this technique is that where the density of original points is higher, the points are moved a shorter distance, resulting in patterns that more closely resemble the original data. The voronoi masking pattern is strongly linked to the residence distribution of the specified study area. Another advantage of voronoi masking is that some points in adjacent polygons will be snapped to the same location, which can increase their k-anonymity.
@@
    Adaptive areal elimination (AAE): This method establishes zones with a minimum level of K-anonymity (Kounadi and Leitner, 2016). Initially, address points are grouped into polygons if available. Subsequently, a disclosure value for ”RoRi” (risk of re-identification) is determined, representing the minimum K-anonymity threshold at which sensitive information can be revealed. The method involves a dissolving process, where polygons with RoRi values below the set threshold are merged with neighbouring polygons until all polygons reach or exceed the threshold. This merging process follows a spatial guideline based on shared borders, primarily prioritising neighbours with the longest shared border. In the case of equilateral polygons, they undergo the dissolving process with all adjacent polygons. Ultimately, this process results in the creation of K-anonymised areas where the disclosure threshold is met. Location points within these areas can be relocated to the centroids of the K-anonymised zones or randomly distributed within them to ensure anonymity.
@@
    Adaptive voronoi masking (AVM): This method integrates the principles of AAE and VM to preserve data utility, considering topological polygon relationships and underlying topography, and achieve a high-level of spatial k-anonymity to mitigate re-identification risks(Polzin, 2020). A key strength of AAE lies in accounting for underlying population density, while VM excels in displacing the Original Data Points (ODP), generating masked patterns similar to the ODP’s distribution. However, both VM and AAE exhibit drawbacks: they overlook the underlying topography, potentially placing data points in illogical locations such as lakes or forests, leading to false re-identification risks. AVM bridges these gaps by integrating the benefits of AAE’s population density consideration and VM’s ODP displacement method. It moves ODPs to the closest segments of their corresponding Voronoi polygon within the merged AAE-polygon, mitigating the risk of displacing points to areas with different population thresholds. Additionally, AVM addresses topographical concerns by relocating data points to the nearest street intersection, minimising false re-identification and illogical placements in water bodies or forests.
@@
    RDV masking: This method, known as the three-layer RDV masking model, is based on rudimentary triangulation and Voronoi generation(Gupta and Rao, 2020). It is intended to be used at the research analysis stage before sharing final data publicly. The masking process is centralised, applying masking to specified nodes at one point. The key advantage of this solution is its ability to not only efficiently mask points but also retain the spatial pattern of the original data, considering various indicators beyond population density. The RDV masking method is structured into three distinctive layers, each contributing to the overall masking process. The first layer, termed layer R, initiates by spatially distributing a given set of points throughout the region, ensuring the preservation of the original point configuration for later comparison and analysis. Moving to layer D, the process progresses to computing the Delaunay triangulation of the given point set, utilising a Divide and Conquer strategy to efficiently accomplish this phase, albeit with a specific computation duration. Finally, layer V, the last phase, focuses on separating individual points to their nearest segment along the edges of the corresponding Voronoi polygon. This step is pivotal in upholding the original density patterns of the points, employing Fortune’s Algorithm for computation, which also entails its specific computation duration. Together, these layers encapsulate a comprehensive approach to masking, offering spatial distribution, triangulation, and precise segmentation to maintain original data patterns effectively.
@@
    Zip4-codes aggregation: This method uses Zip4 (postal) codes as a spatial aggregation unit for sharing fine-scale health data while addressing privacy concerns (Ajayakumar et al., 2023). The Zip4 aggregation preserves the underlying spatial structure making it potentially suitable for geospatial analysis.
@@ 3.1.4.3 Differentially private clustering
Differential privacy is the current state-of-the-art technique for controlling risk of identity disclosure, that greatly reduces the disclosure risk issues in k-anonymity, l-diversity and other extensions by leveraging randomisation with noise addition (Dwork, 2006). An algorithm is said to be differentially private if the operation of that algorithm on the data set produces essentially the same answers regardless of the presence or absence of an individual.
@@
The main drawback of using differentially private algorithms for data analysis is that, like the other perturbation methods, it does not preserve the accuracy of the algorithms applied to the original data, and in fact will deviate more from the results based on the original data in proportion to the privacy guarantees that it provides (Dwork, 2008). Further, differential privacy does implicitly assume that all information about an individual participant is contained within a record and that records are independent. It is not appropriate to input a data set containing multiple records per person, into a differentially-private mechanism given if records are correlated, the privacy protections may become weaker.
Differential privacy has been used for querying geospatial data sets in a privacy preserving manner (Xiao and Xiong, 2015; Harris, 2020). Apart from applying differential privacy for statistical queries, differential privacy has been utilised for geospatial clustering. Clustering on geospatial data can inadvertently create vulnerabilities that adversaries with background knowledge can exploit to breach individuals’ privacy. To safeguard individual privacy from such risks, significant research has been devoted to preserving differential privacy in clustering algorithms (Yan et al., 2022; Wang et al., 2017; Qardaji et al., 2013; Lu and Shen, 2020). Figure 7 illustrates how original k-means clustering (left plot) and how differential privacy based k-means clustering (right plot) protects location privacy.
@@ 3.1.5. Synthetic data generation
This technique allows mixing up the elements of an dataset or creating new values based on the original information—so the overall totals, values and patterns of the data are preserved but do not relate to any particular individual (Sakshaug and Raghunathan, 2014). Synthetic data can be generated in a number of ways. (1) Using a generative computer model or generated adversarial network (GAN) to create a set of data points that cannot be differentiated from the real data. In generative models, algorithms are fed with smaller real-world data which then gets derived by the algorithms to create similar data. (2) Enhanced sampling either with over-sampling the minority class or under-sampling the majority case to create a synthetic distribution of data. Finally, (3) an agent based simulation which uses a simulation process where agents are developed to represent real-world entities that interact with each other, and these interactions are observed and measured to generate synthetic data.
Using above methods, the data can be generated either partially or fully synthetic (Hu and Drechsler, 2015). In partially synthetic data only attribute values that are sensitive are replaced with synthetic data. In fully synthetic data generation an entire data set (rather than a sample of data) is replaced with synthesised replacement data. However, synthetic data is dependent on the model that is used to generate or derive synthesised replacement data. If the initial model is not sufficient to create quality synthetic data, the generated data will be less reflective of original data.
Synthetic data generation becomes especially crucial when working with geocoded data, such as addresses or coordinates, as releasing them in their raw format can introduce substantial privacy risks (Nowok and Dibben, 2018). Synthetic data that replicates the structure and statistical properties of the original geospatial data set without using original content offer a way to enhance the use of confidential microdata, also with detailed geographic identifiers. These data sets can be used to conduct reliable inference in many analyses but in some cases they will give approximate results at best. The choice of a specific synthetic data set for release from all the versions created depends on the intended level of geographical analysis and the desirable risk-utility profile (Quick et al., 2015).
@@ 3.2. Evaluation measure for Geomasking
The performance of the geomasking techniques can be evaluated in terms of how efficient and effective they are when applied on geospatial data sets. The efficiency of a geomasking technique evaluate how scalable a masking technique is on a large data set with millions of records, while the effectiveness of a geomasking technique is measured by the accuracy and the protection it can provide for an individual. Next, we discuss different measures that have been proposed to evaluate geomasking techniques in terms of scalability, accuracy, and privacy.
@@ 3.2.1. Scalability
Scalability can be evaluated using measures that are dependent on the computing platform and networking infrastructure used, or measures that are based on the number of candidate record pairs generated. Common measures include run time, memory space, and communication size.
Apart from these measures, the scalability can also be evaluated theoretically using computational complexity measures. This includes the computation and communication complexities (if multiple data providers are available) that measure the overall computational efforts and cost of communication required in the geomasking process. Generally, the big O-notation is used to specify the computation complexity. Given $n$ is the number of records in a database, the big O-notation of represents $O(log\ n)$ logarithmic complexity, $O(n)$ linear complexity, $O(n\cdot log\ n)$ log-linear complexity, $O(n^{2})$ quadratic complexity, $O(n^{c})$ polynomial complexity, $O(polylog\ n)$ polynomial logarithmic complexity, and exponential complexity, where $c>1$.
@@ 3.2.2. Accuracy
We classify the metrics used for evaluating the accuracy of geomasking techniques into Spatial Point Pattern & Distribution Analysis, Displacement Distance Analysis, and Spatial Autocorrelation Analysis.
@@ 3.2.2.1 Spatial Point Pattern & Distribution Analysis
    Average Nearest Neighbour (ANN): ANN is a spatial analysis method used to quantify the arrangement of points within a dataset(Iyer et al., 2023; Wang et al., 2022; Ajayakumar et al., 2023; Broen et al., 2021; Seidl et al., 2015; Zandbergen, 2014). It calculates the average distance of each point to its nearest neighbour, offering valuable insights into the spatial pattern of the data. This metric is particularly useful for identifying clustering at a global level providing valuable information about the spatial distribution and clustering tendencies of the dataset.
@@ 3.2.3. Privacy
In this section we discuss several metrics that can be used to measure the privacy of the geomasking techniques.
@@
Apart from the above measures, Redlich (Redlich, 2022) highlighted additional re-identification risk assessments. These measures include linking the geodata with a dataset containing longitudinal information, mean of multiple release, and reversing masking techniques. Linking geodata with longitudinal information involves mapping masked geo-spatial records with a dataset containing direct identifiers about individuals, such as names and addresses.
For the linking process, two attack schemes have been proposed. (1) Graph theoretic linkage attack (Kroll, 2015) uses graph theory to link masked and original coordinates, framing the problem as a maximum clique issue. (2) Graph matching attack (Vidanage et al., 2020) uses graph matching mechanisms to identify matching records in an encoded and plaintext datasets. The attack utilises features like node-based, edge-based, and structural features, applying cosine locality-sensitive hashing for identifying matching records. The second measure, mean of multiple releases, involves using various dataset versions to determine the original coordinates’ location. The idea is based on the central limit theorem, which states that as the number of observations increases, the average of the masked coordinates will fall closer to the original point . The third method reverse masking methods aim to reverse a masking method, such as affine transformation, by testing all possible parameters to find the original location.
@@ 3.3. Practical aspects of Geomasking
The final three dimensions cover practical aspects of geomasking techniques including the datasets used for experimental evaluations, how a solution was implemented, and if a proposed solution was developed with a specific application area in mind.
@@ 3.3.1. Data sets
To evaluate the practical aspects of geomasking techniques with regard to their expected performance in real-world applications, evaluations should ideally be done on databases that exhibit real-world properties and error characteristics. However, in some occasions, due to the difficulties of obtaining real-world data that contain geospatial information, synthetically generated geo databases could also be used. Table 2 provides details of several datasets that have been used by different studies for experiments with geomasking techniques. It is important to note that though these datasets are available publicly, some of these datasets might have ethical obligations.
Table 2. Datasets used by different studies for experiments on geo-masking techniques.
@@
European data
The official portal for European data
@@
https://data.europa.eu/data/datasets/?locale=en
(Polzin, 2020)
French public data
@@
Open platform for French public data. (Houfaf-Khoufaf et al., 2021) utilised below datasets - COVID-19 virological test results on https://www.data.gouv.fr/fr/datasets/old-donnees-relatives-aux-resultats-des-tests-virologiques-covid-19/ - Base National Address on https://www.data.gouv.fr/fr/datasets/base-adresse-nationale/
@@
https://www.data.gouv.fr/fr/datasets/
(Houfaf-Khoufaf et al., 2021)
US Census Bureau data
@@
This repository provides data related to People, Places, and Economy; covering over 100,000 different geographies: States, counties, places, tribal areas, zip codes, and congressional districts in the United States. For each, they cover topics, including but not limited to education, employment, health, and housing. - (Wang et al., 2022) utilised United States Census Bureau, 2021 - (Dupre et al., 2022) utilised United States Census Bureau, 2019 (Income and Poverty in the United States: 2019)
https://data.census.gov/
(Wang et al., 2022; Dupre et al., 2022)
US GIS data
@@
The NYS GIS Clearinghouse is an evolving searchable repository of GIS data and mapping resources. (Zhang et al., 2017) utilised actual residential datasets from three U.S. counties in their analysis, specifically examining data from Jackson County in Oregon, Travis County in Texas, and Wake County in North Carolina.
https://data.gis.ny.gov/
(Zhang et al., 2017)
U.S. Census Bureau’s geographic spatial data
@@
This repository represents the U.S. Census Bureau’s geographic spatial data which has been developed using Topologically Integrated Geographic Encoding and Referencing system (TIGER) geospatial data as the primary source.
https://www.census.gov/programs-surveys/geography/guidance/tiger-data-products-guide.html
(Su et al., 2017)
Australian government open data
@@
Data.gov.au serves as the primary repository for open government data in Australia. Individuals can retrieve anonymised public data released by federal, state, and local government agencies.
https://data.gov.au/search
(Redlich, 2022)
Geoscape Geocoded National Address File (G-NAF)
@@
Geoscape G-NAF is a trusted source of geocoded address database for both Australian businesses and governmental entities, comprising more than 50 million submitted addresses condensed into 15.4 million G-NAF addresses. It is built and maintained by Geoscape Australia using independently examined and validated government data. (Redlich, 2022) utilised August 2019 - PSMA Geocoded National Address File (G-NAF) from https://data.gov.au/dataset/ds-dga-e1a365fc-52f5-4798-8f0c-ed1d33d43b6d/distribution/dist-dga-32be073c-338b-40df-9842-7cd91a25d960/details?q=
https://data.gov.au/dataset/ds-dga-19432f89-dc3a-4ef3-b943-5326ef1dbecc/details
(Redlich, 2022)
Open platform for Research data (OSF)
OSF serves as a freely accessible and open platform designed to facilitate research and foster collaboration by providing access to projects, data, and materials that could be beneficial to the research process. - https://osf.io/6uqkv: provided data by (Swanlund et al., 2020b) - https://osf.io/hxd5w/: provided data by (Wang et al., 2022)
https://osf.io/search?resourceType=File
(Swanlund et al., 2020b; Wang et al., 2022)
Earth Engine Data Catalogue
@@
The public data catalogue of Earth Engine encompasses a range of standard raster datasets related to Earth science. (Clarke, 2016) utilised a public GNSS track from a hike on the Tennessee Valley trail in the Golden Gate National Recreation Area, California, retrieved from the user-contributed data for Google Earth. The track consists of 360 points specified in geographic decimal degrees on the WGS84 coordinate system.
https://developers.google.com/earth-engine/datasets/catalog
(Clarke, 2016)
Deaths recorded in Lawrence, Massachusetts from 1911 to 1913
@@
This repository, provided by (Broen et al., 2021), contains geocoded data on all deaths that occurred in Lawrence, MA, from 1911 to 1913. Additionally, it provides the code for applying various spatial perturbations and generating descriptive spatial statistics.
https://github.com/broenk/Spatial_Perturbation
(Broen et al., 2021)
GIS data for the city of Lawrence, Massachusetts
@@
This repository provides GIS data for the city of Lawrence, Massachusetts.
https://data.mass.gov/browse/gis
(Broen et al., 2021)
GIS data for Orange County, North Carolina
@@
This repository provides GIS data for Orange County, North Carolina. (Allshouse et al., 2010) employed the E911 database for Orange County, North Carolina (Orange County GIS Division 2007), which includes census block groups exhibiting diverse population densities.
https://www.orangecountync.gov/1610/Land-Records-GIS-Services
(Allshouse et al., 2010)
Bike sharing dataset (NY, USA)
@@
This repository provides citi bike trip data, including details on their riding locations, preferred timing, ride distances, popular stations, and the days of the week with the highest ride frequency.
https://citibikenyc.com/system-data
(Hasan et al., 2017)
Austrian statistical data
@@
The platform data.statistik.gv.at provides various datasets sourced from the official Austrian statistical data dictionary. These datasets adhere to the Principles Of Open Data, ensuring their accessibility in machine-readable formats. In the study by (Seidl et al., 2015), energy consumption data for each household in the Hermagor district of Carinthia in southern Austria was utilised. This dataset encompasses 1945 residential records, represented by the centroids of georeferenced buildings supplied by the respective communities in the district.
https://www.statistik.at/datenbanken/statcube-statistische-datenbank
(Seidl et al., 2015)

### v2

@@ 3.1.4.2 Spatial tessellation-based techniques
    Voronoi masking: Voronoi polygons define areas where the boundaries are equidistant between surrounding points or the interior of the polygons is closer to the corresponding point than to any other points (Seidl et al., 2015). In Voronoi masking, each point is snapped to the closest part along the edges of its corresponding Voronoi polygon. An advantage of this technique is that where the density of original points is higher, the points are moved a shorter distance, resulting in patterns that more closely resemble the original data. The Voronoi masking pattern is strongly linked to the distribution of points within the specified study area. Another advantage of Voronoi masking is that some points in adjacent polygons will be snapped to the same location, which can increase their k-anonymity.
@@
    Adaptive areal elimination (AAE): This method establishes zones with a minimum level of K-anonymity (Kounadi and Leitner, 2016). Initially, data points are grouped into polygons if available. Subsequently, a disclosure value for ”RoRi” (risk of re-identification) is determined, representing the minimum K-anonymity threshold at which sensitive information can be revealed. The method involves a dissolving process, where polygons with RoRi values below the set threshold are merged with neighbouring polygons until all polygons reach or exceed the threshold. This merging process follows a spatial guideline based on shared borders, primarily prioritising neighbours with the longest shared border. In the case of equilateral polygons, they undergo the dissolving process with all adjacent polygons. This process results in the creation of K-anonymised areas where the disclosure threshold is met. Location points within these areas can be relocated to the centroids or randomly distributed within the K-anonymised zones.
@@
    Adaptive voronoi masking (AVM): This method integrates the principles of AAE and VM to preserve data utility, considering topological polygon relationships and underlying topography, and achieve a high level of spatial k-anonymity to mitigate re-identification risks(Polzin, 2020). A key strength of AAE lies in accounting for underlying population density, while VM excels in displacing the Original Data Points (ODP), generating masked patterns similar to the ODP’s distribution. However, both VM and AAE exhibit drawbacks: they overlook the underlying topography, potentially placing data points in illogical locations such as lakes or forests, leading to false re-identification risks. AVM bridges these gaps by integrating the benefits of AAE’s population density consideration and VM’s ODP displacement method. It moves ODPs to the closest segments of their corresponding Voronoi polygon within the merged AAE-polygon, mitigating the risk of displacing points to areas with different population thresholds. Additionally, AVM addresses topographical concerns by relocating data points to the nearest street intersection, minimising false re-identification and illogical placements.
@@
    RDV masking: This method, known as the three-layer RDV masking model, is based on rudimentary triangulation and Voronoi generation(Gupta and Rao, 2020). The masking process is centralised, applying masking to specified nodes at one point. The key advantage of this solution is its ability to not only efficiently mask points but also retain the spatial pattern of the original data, considering various indicators beyond population density. The RDV masking method is structured into three distinctive layers, each contributing to the overall masking process. The first layer, termed layer R, initiates by spatially distributing a given set of points throughout the region, ensuring the preservation of the original point configuration for later comparison and analysis. Moving to layer D, the process progresses to computing the Delaunay triangulation of the given point set, utilising a Divide and Conquer strategy to efficiently accomplish this phase, albeit with a specific computation duration. Finally, layer V, the last phase, focuses on separating individual points to their nearest segment along the edges of the corresponding Voronoi polygon. This step is pivotal in upholding the original density patterns of the points, employing Fortune’s Algorithm for computation, which also entails its specific computation duration. Together, these layers encapsulate a comprehensive approach to masking, offering spatial distribution, triangulation, and precise segmentation to maintain original data patterns effectively.
@@
    Zip4codes aggregation: This method uses Zip4 (postal) codes as a spatial aggregation unit for sharing fine-scale health data while addressing privacy concerns (Ajayakumar et al., 2023). The Zip4 aggregation preserves the underlying spatial structure, making it potentially suitable for geospatial analysis.
@@ 3.1.4.3 Differentially private clustering
Differential privacy (DP), recognised as the current state-of-the-art provable privacy technique for mitigating the risk of identity disclosure by incorporating randomisation and noise addition (Dwork, 2006). By leveraging randomisation techniques coupled with noise addition, DP significantly diminishes the disclosure risk associated with these methods, offering robust protection for sensitive data. An algorithm is said to be differentially private if the algorithm produces essentially the same answers regardless of the presence or absence of an individual in the dataset.
@@
DP has recently been used for querying geospatial datasets in a privacy-preserving manner (Xiao and Xiong, 2015; Harris, 2020). In addition to its application in statistical queries, DP has been employed in geospatial clustering. Clustering geospatial data can inadvertently expose vulnerabilities. For example, the disclosure of cluster centroids during the iterative clustering process can lead to the potential leakage of information about individual data points, particularly when adversaries have background knowledge or auxiliary information enabling them to conduct inference attacks (Lu and Shen, 2020).
DP techniques could inject noise into the clustering process such that the centroids may move slightly or significantly from their original positions depending on the magnitude and distribution of the injected noise (Yan et al., 2022; Wang et al., 2017; Qardaji et al., 2013; Lu and Shen, 2020). Figure 7 illustrates both the original k-means clustering (on the left) and the differential privacy-based k-means clustering (on the right) applied to the same dataset. To mitigate such risks and safeguard individual privacy, considerable research effort has been dedicated to integrating differential privacy into clustering algorithms.
The main drawback of using DP algorithms for data analysis is that, like other perturbation methods, they do not preserve the accuracy of the algorithms applied to the original data. In fact, they tend to deviate more from the results based on the original data in proportion to the privacy guarantees they provide (Balcan et al., 2017). Further, DP implicitly assumes that all information about an individual participant is contained within a record and that records are independent. Thus, if the dataset contains multiple location records per person the the DP mechanism needs to be design to capture all these records of an individual. If location records are correlated, the privacy protections may weaken, as the DP mechanism may inadvertently leak information due to the interdependence of records.
@@ 3.1.5. Synthetic data generation
Synthetic datasets aim to preserve the spatial characteristics and patterns present in the original data while preserving the individual’s privacy. Synthetic data can be generated in a number of ways (Bauer et al., 2024). Some synthetic data generation techniques include (1) Generative AI; these techniques are based on Generative Pre-trained Transformers (GPT), Generative Adversarial Networks (GANs), or Variational Auto-Encoders (VAEs) which learn the underlying distribution of real data to generate similarly distributed synthetic data (Zhao et al., 2021; Wu and Biljecki, 2022). (2) A rules engine that can create synthetic data via user-defined business rules. Intelligence can be added to the generated data by referencing the relationships between the data elements, to ensure the relational integrity of the generated data (McKenzie et al., 2023). (3) Entity cloning is another technique that extracts data from the source systems of a single entity (e.g., patient) and masks it for compliance. It then clones the entity, generating different identifiers for each clone to ensure uniqueness. (4) Data masking is also a possible option that replaces personally identifiable information with fictitious, yet structurally consistent, values, such as addresses or coordinates (Nowok and Dibben, 2018). The objective of data masking is to ensure that sensitive data cannot be linked to individuals, while retaining the overall relationships and statistical characteristics of the data.
Some other techniques, such as Copula models, could discover correlations and dependencies within the datasets, and then use these correlations and dependencies to generate new, realistic data. Data augmentation (such as flipping, rotation, scaling, and translation) could also be applied as supplementary techniques to existing values to create new data. Finally, Techniques such as random sampling and noise injection, which add data points from known distributions, could also be used to create new data points that closely resemble real-world data (Sakshaug and Raghunathan, 2014).
@@ 3.2. Evaluation measure for Geomasking
The performance of the geomasking techniques can be evaluated in terms of how efficient and effective they are when applied to geospatial datasets. The efficiency of a geomasking technique evaluates how scalable a masking technique is on a large dataset with millions of records, while the effectiveness of a geomasking technique is measured by the accuracy and the protection it can provide for an individual. Next, we discuss different measures that have been proposed to evaluate geomasking techniques in terms of scalability, accuracy, and privacy.
@@ 3.2.1. Scalability
Scalability can be evaluated using measures that are dependent on the computing platform and networking infrastructure used, or measures that are based on the number of data points analysed (Redlich, 2022). Common measures include run time, memory space, and communication size.
Apart from these measures, the scalability can also be evaluated theoretically using computational complexity measures. This includes the computation and communication complexities (if multiple data providers are available) that measure the overall computational efforts and cost of communication required in the geomasking process. Generally, the big O-notation is used to specify the computation complexity (Papadimitriou, 2003). Given $n$ is the number of records in a database, the big O-notation of represents $O(log\ n)$ logarithmic complexity, $O(n)$ linear complexity, $O(n\cdot log\ n)$ log-linear complexity, $O(n^{2})$ quadratic complexity, $O(n^{c})$ polynomial complexity, $O(polylog\ n)$ polynomial logarithmic complexity, and exponential complexity, where $c>1$.
@@ 3.2.2. Accuracy
Table 1 presents our classification of metrics for evaluating the accuracy of geomasking techniques. These metrics are categorised into three groups: Spatial Point Pattern & Distribution Analysis, Displacement Distance Analysis, and Spatial Autocorrelation Analysis.
@@ 3.2.2.1 Spatial Point Pattern & Distribution Analysis
    Average Nearest Neighbour (ANN): ANN is a spatial analysis method used to quantify the arrangement of points within a dataset(Iyer et al., 2023; Wang et al., 2022; Ajayakumar et al., 2023; Broen et al., 2021; Seidl et al., 2015; Zandbergen, 2014). It calculates the average distance of each point to its nearest neighbour, offering valuable insights into the spatial pattern of the data. This metric is particularly useful for identifying clustering at a global level, providing valuable information about the spatial distribution and clustering tendencies of the dataset.
@@ 3.2.3. Privacy
In this section, we discuss several metrics that can be used to measure the privacy of the geomasking techniques.
@@
Apart from the above measures, Redlich (Redlich, 2022) highlights additional re-identification risk assessments. These measures include linking the geodata with a dataset containing longitudinal information, mean of multiple releases, and reversing masking techniques. Linking geodata with longitudinal information involves mapping masked geospatial records with a dataset containing direct identifiers about individuals, such as names and addresses.
For the linking process, two attack schemes have been proposed. (1) Graph theoretic linkage attack (Kroll, 2015) uses graph theory to link masked and original coordinates, framing the problem as a maximum clique issue. (2) Graph matching attack (Vidanage et al., 2020) uses graph matching mechanisms to identify matching records in encoded and plaintext datasets. The attack utilises features like node-based, edge-based, and structural features, applying cosine locality-sensitive hashing for identifying matching records. The second measure, mean of multiple releases, involves using various dataset versions to determine the original coordinates’ location. The idea is based on the central limit theorem, which states that as the number of observations increases, the average of the masked coordinates will fall closer to the original point. The third measure, reversing masking techniques, aims to reverse a masking method, such as affine transformation, by testing all possible parameters to find the original location.
@@ 3.3. Practical aspects of Geomasking
The final three dimensions cover practical aspects of geomasking techniques, including the datasets used for experimental evaluations, how a solution was implemented, and various application domains where geomasking can play a pivotal role.
@@ 3.3.1. Datasets
To evaluate the practical aspects of geomasking techniques with regard to their expected performance in real-world applications, evaluations should ideally be done on databases that exhibit real-world properties and error characteristics. However, in some occasions, due to the difficulties of obtaining real-world data that contain geospatial information, synthetically generated geo databases could also be used. Table LABEL:tab:datasets provides details of several datasets that have been used by different studies for experiments with geomasking techniques. It is important to note that though these datasets are available publicly, some of these datasets might have ethical obligations.
Table 2. Datasets used by different studies for experiments on geomasking techniques.
@@
European data The official portal for European data https://data.europa.eu/data/datasets/?locale=en (Polzin, 2020)
@@
French public data Open platform for French public data. (Houfaf-Khoufaf et al., 2021) utilised below datasets - COVID-19 virological test results on
@@
https://www.data.gouv.fr/fr/datasets/old-donnees-relatives-aux-resultats-des-tests-virologiques-covid-19/
- Base National Address on https://www.data.gouv.fr/fr/datasets/base-adresse-nationale/
https://www.data.gouv.fr/fr/datasets/ (Houfaf-Khoufaf et al., 2021)
@@
US Census Bureau data This repository provides data related to People, Places, and Economy; covering over 100,000 different geographies: States, counties, places, tribal areas, zip codes, and congressional districts in the United States. For each, they cover topics including, but not limited to, education, employment, health, and housing. - (Wang et al., 2022) utilised United States Census Bureau, 2021 - (Dupre et al., 2022) utilised United States Census Bureau, 2019 (Income and Poverty in the United States: 2019) https://data.census.gov/ (Wang et al., 2022; Dupre et al., 2022)
@@
US GIS data The NYS GIS Clearinghouse is an evolving searchable repository of GIS data and mapping resources. (Zhang et al., 2017) utilised actual residential datasets from three U.S. counties in their analysis, specifically examining data from Jackson County in Oregon, Travis County in Texas, and Wake County in North Carolina. https://data.gis.ny.gov/ (Zhang et al., 2017)
@@
U.S. Census Bureau’s geographic spatial data This repository represents the U.S. Census Bureau’s geographic spatial data, which has been developed using Topologically Integrated Geographic Encoding and Referencing system (TIGER) geospatial data as the primary source. https://www.census.gov/programs-surveys/geography/guidance/tiger-data-products-guide.html (Su et al., 2017)
US GIS data The NYS GIS Clearinghouse is an evolving searchable repository of GIS data and mapping resources. (Zhang et al., 2017) utilised actual residential datasets from three U.S. counties in their analysis, specifically examining data from Jackson County in Oregon, Travis County in Texas, and Wake County in North Carolina. https://data.gis.ny.gov/ (Zhang et al., 2017)
U.S. Department of Transportation (USDOT) the National Address Database (NAD), established by the USDOT and its partners, provides addresses aggregated from state, local, and tribal governments. These addresses are compiled to provide accurate and up-to-date information critical for various purposes, including transportation safety, Next Generation 9-1-1 services, mail delivery, permitting, and school siting. https://www.transportation.gov/gis/national-address-database (Lin, 2023)
@@
Australian Government open data Data.gov.au serves as the primary repository for open government data in Australia. Individuals can retrieve anonymised public data released by federal, state, and local government agencies. https://data.gov.au/search (Redlich, 2022)
@@
Geoscape Geocoded National Address File (G-NAF) Geoscape G-NAF is a trusted source of geocoded address database for both Australian businesses and governmental entities, comprising more than 50 million submitted addresses condensed into 15.4 million G-NAF addresses. It is built and maintained by Geoscape Australia using independently examined and validated government data. (Redlich, 2022) utilised August 2019 - PSMA Geocoded National Address File (G-NAF) from https://data.gov.au/dataset/ds-dga-e1a365fc-52f5-4798-8f0c-ed1d33d43b6d/distribution/dist-dga-32be073c-338b-40df-9842-7cd91a25d960/details?q= https://data.gov.au/dataset/ds-dga-19432f89-dc3a-4ef3-b943-5326ef1dbecc/details (Redlich, 2022)
@@
Earth Engine Data Catalogue The public data catalogue of Earth Engine encompasses a range of standard raster datasets related to Earth science. (Clarke, 2016) utilised a public GNSS track from a hike on the Tennessee Valley trail in the Golden Gate National Recreation Area, California, retrieved from the user-contributed data for Google Earth. The track consists of 360 points specified in geographic decimal degrees on the WGS84 coordinate system. https://developers.google.com/earth-engine/datasets/catalog (Clarke, 2016)
@@
Deaths recorded in Lawrence, Massachusetts from 1911 to 1913 This repository, provided by (Broen et al., 2021), contains geocoded data on all deaths that occurred in Lawrence, MA, from 1911 to 1913. Additionally, it provides the code for applying various spatial perturbations and generating descriptive spatial statistics. https://github.com/broenk/Spatial_Perturbation (Broen et al., 2021)
@@
GIS data for the city of Lawrence, Massachusetts This repository provides GIS data for the city of Lawrence, Massachusetts. https://data.mass.gov/browse/gis (Broen et al., 2021)
@@
GIS data for Orange County, North Carolina This repository provides GIS data for Orange County, North Carolina. (Allshouse et al., 2010) employed the E911 database for Orange County, North Carolina (Orange County GIS Division 2007), which includes census block groups exhibiting diverse population densities. https://www.orangecountync.gov/1610/Land-Records-GIS-Services (Allshouse et al., 2010)
@@
Bike sharing dataset (NY, USA) This repository provides citi bike trip data, including details on their riding locations, preferred timing, ride distances, popular stations, and the days of the week with the highest ride frequency. https://citibikenyc.com/system-data (Hasan et al., 2017)
@@
Austrian statistical data The platform data.statistik.gv.at provides various datasets sourced from the official Austrian statistical data dictionary. These datasets adhere to the Principles Of Open Data, ensuring their accessibility in machine-readable formats. In the study by (Seidl et al., 2015), energy consumption data for each household in the Hermagor district of Carinthia in southern Austria was utilised. This dataset encompasses 1945 residential records, represented by the centroids of georeferenced buildings supplied by the respective communities in the district. https://www.statistik.at/datenbanken/statcube-statistische-datenbank (Seidl et al., 2015)
Open platform for research data (OSF) OSF serves as a freely accessible and open platform designed to facilitate research and foster collaboration by providing access to projects, data, and materials that could be beneficial to the research process. - Synthetic clusters, addresses, and masked data provided by (Swanlund et al., 2020b): https://osf.io/6uqkv: - Simulated data provided by (Wang et al., 2022): https://osf.io/hxd5w/ https://osf.io/search?resourceType=File (Swanlund et al., 2020b; Wang et al., 2022)
Figshare Figshare is an open cloud-based platform used for managing, sharing, and discovering research outputs, including data, figures, code, presentations, and other types of scholarly content. - A synthetic address-level dataset provided by (Lin, 2023): https://figshare.com/articles/dataset/Geo-indistinguishable_masking_Enhancing_privacy_protection_in_spatial_point_mapping/23632443 https://osf.io/search?resourceType=File (Lin, 2023)
3.3.2. Implementations
This dimension specifies the applications and prototypes that have been developed and implemented for various geoprivacy techniques in order to conduct experimental evaluation. Some solutions proposed in the literature provide only algorithmic details, but they have not been evaluated experimentally, or no details about their implementation have been published. Table 3 provides a brief description of existing tools implementing different geomasking techniques.